OpenStudy (anonymous):

Could someone help me solve these problems: 1. |-3i| 2. (-1 + 2i) + (6-9i) 3. 3+2i/4+ i PLEASE HELP!!!

OpenStudy (anonymous):

1. 3i

OpenStudy (anonymous):

What about 2 and 3?

jimthompson5910 (jim_thompson5910):

hint: |a+bi| = sqrt(a^2 + b^2)

OpenStudy (anonymous):

Yaa could you break it down i am new to this!

jimthompson5910 (jim_thompson5910):

-3i = 0-3i = 0 + (-3i)

jimthompson5910 (jim_thompson5910):

so -3i is in the form a+bi where a = 0 and b = -3

jimthompson5910 (jim_thompson5910):

plug in a = 0 and b = -3 into |a+bi| = sqrt(a^2 + b^2)

OpenStudy (anonymous):

ok thanks i think i understand it know!

jimthompson5910 (jim_thompson5910):

you're welcome